OpenStudy (anonymous):

simplify x2-2x-3 over x2-7x+12

OpenStudy (anonymous):

\[\frac{x^2-2x-3}{x^2-7x+12}=\frac{(x-3)(x+1)}{(x-3)(x-4)}=\frac{x+1}{x+4}\]